# A Tidy Data Structure and Visualisations for Multiple Variable Correlations and Other Pairwise Scores

### v1

@@ 4   Bullseye visualisations
Next, we investigate the associations for all variables grouped by species. The code below uses pairwise_scores to calculate pairwise associations grouped by species. We use var_order="seriate_max_diff" to bring variable pairs with the biggest score difference to the top-left corner. Figure 2 shows the result.
sc_species<- pairwise_scores(penguins, by="species")
plot(sc_species, var_order="seriate_max_diff")
@@
Figure 2: Pairwise correlations for numerical variables and canonical correlation for factors and mixed factor-numeric pairs grouped by species. Ungrouped measures are in the centre bullseye, outer wedges represent Adelie, Chinstrap, Gentoo in anti-clockwise order. The grey colour represents NA values. Tooltips show the groups and the score values in the html version. A blue inner circle surrounded by red wedges indicates that direction of correlation is reversed within groups for some variable pairs.
@@
The inner circle of Figure 2 shows the ungrouped score, and the outer doughnut shows the scores for each of the species. The html version of this plot is interactive, which is useful to obtain the species represented in each segment and the numerical value of the score. Generally, it is of interest if there is a high difference among the grouped and ungrouped scores. Here the variable bill_dep has a negative overall correlation with each of flip_len, body_mass, bill_len but positive correlation within each of the species groups. The same pattern (known as Simpson’s paradox) is present for the pair bill_len and bill_dep. There are also a number of wedges coloured in grey involving the variable island, representing score values of NA. This is because two of the species (Gentoo and Chinstrap) are located on one island apiece only. For these species, island is constant and so the association value cannot be calculated and is therefore recorded as NA.
@@
To demonstrate the linear display available in bullseye, we switch to a dataset with more variables. We use the acs12 dataset from the package openintro (Çetinkaya-Rundel et al. 2024) which contains results from the US Census American Community Survey in 2012. The dataset has 13 variables, only four of which (income, hrs_work, age, time_to_work) are numeric, the rest are factors. The code segment below calculates ace, cancor and nmi scores. (These three measures were chosen because they are applicable to both numerical and factor variables.)
@@
A display similar to Figure 2 without the center segments could be constructed with the code
plot(sc_multi, interactive=TRUE)
However, rather than displaying all (${{13}\choose{2}}=78$ variable pairs we chose instead to filter out the interesting variable pairs from sc_multi. In this example, we show variable pairs whose maximum score across the three measures exceeds 0.25. The filtered pairwise structure no longer has all (x,y) pairs, so we use the linear display as shown in Figure 3 instead of the matrix display.
@@
  plot(type="linear")
@@
In Figure 3 variable pairs are sorted from highest to lowest maximum score. We see that the mutual information measure (nmi) is consistently less than the ace score. By definition the ace score cannot be less than the cancor score, and they are identical for categorical pairs of variables, for example, (citizen, lang) and (lang, race). The ace scores for income with other variables is much higher than the cancor score, indicating non-linear association. For example, an investigation of the (age, income) scatterplot shows that income goes up with age until about age 40 and then drops off.
@@
group_scores <- pairwise_scores(acs12, by = "race")
rng <- function(vals){
  if (all(is.na(vals))) 0 else max(vals, na.rm=TRUE)- min(vals,na.rm=TRUE)
mutate(group_scores, valrange = rng(value),valmax = max(abs(value)), .by=c(x,y))|>
@@
  plot(type="linear", geom="point", pair_order = "seriate_max_diff")+
   theme(legend.text = element_text(size = rel(.5)),
         legend.title = element_text(size = rel(.5))
@@ 5   Bullseye and other packages
Finally, the pairwise structure is easily manipulated to construct general plots: The example code below and resulting display in Figure 5 shows that pivoting the sc_multi pairwise structure containing scores calculated on the acs12 data, gives a dataset where the relationship between the scores can easily be compared using a standard ggplot2 visualisation. Here we see that for these data there is an increasing trend in the relationship between ace and nmi scores, though aces scores are higher.
@@ 6   Discussion
In this paper we described a new R package bullseye which fills a gap in providing a standard interface for the calculation of (grouped) pairwise variable scores, a tidy data structure for their manipulation, and a new, optionally interactive, visualisation for the display of multiple pairwise scores. As demonstrated in Section 4 the visualisation is particularly useful for easy identification of the presence of Simpson’s paradox. A second linear display is useful to hone in on a filtered set of interesting pairwise scores.
As explained in Section 3, the pairwise constructor assumes that pairwise variable scores are symmetric, and similarly at present the visualisation functions assume symmetry. It would however be straightforward to extend to asymmetric measures and visualisations, and furthermore to include scores and visualisations for single variables.
Of course, there are a plethora of methods for measuring pairwise patterns which are not included in our package. For example, Tjøstheim, Otneim, and Støve (2022) provides a survey of recent developments in measures of dependence for numerical variables. As discussed in Section 4.3, bullseye visualisations are easily accessible for pairwise scores implemented in R beyond those implemented in our package.
The interactive capabilities of bullseye are currently limited to providing informative tooltips. In future, our package could be used as vehicle for interactively displaying interesting pairwise variable plots. Possibilities in this area have been explored by Grimm (2016), who in the archived package mbgraphic, used correlation displays and scagnostics to identify informative scatterplot matrices and parallel coordinate plots embedded in an interactive Shiny application, and also by Dang and Wilkinson (2014), though their work is not implemented in an R package.
@@ 7   Acknowledgement
This publication has emanated from research supported by a grant from Science Foundation Ireland under Grant number 18/CRT/6049. For the purpose of Open Access, the authors have applied a CC BY public copyright licence to any Author Accepted Manuscript version arising from this submission.

### v2

@@ 4  Bullseye functionality
Next, we investigate the associations for all variables grouped by species. The code below uses pairwise_scores to calculate pairwise associations grouped by species. Figure 3 shows the result.
pairwise_scores(penguins, by="species",add.nobs = TRUE) |>
  plot(var_order="seriate_max_diff")
@@
Figure 3: Pairwise correlations for numerical variables and canonical correlation for factors and mixed factor-numeric pairs grouped by species for the penguin dataset. Ungrouped measures are in the centre bullseye, outer wedges represent Adelie, Chinstrap, Gentoo in anti-clockwise order. The grey colour represents NA values. Tooltips show the groups and the score values in the html version. A blue inner circle surrounded by red wedges indicates that direction of correlation is reversed within groups for some variable pairs.
Note that the plot_pairwise function has an argument var_order which specifies the seriation algorithm to be used to arrange the variables. The default is var_order = "seriate_max" which brings variable pairs with the biggest score value to the top-left corner. This was used to construct both Figures 1 and 2. Alternatively, a value of var_order = "seriate_max_diff" brings variable pairs with the biggest score difference to the top-left corner, which is a useful strategy to uncover interesting group-wise patterns, as shown in Figure 3. Another option is to supply var_order as a vector of variable names.
@@
The inner circle of Figure 3 shows the ungrouped score, and the outer doughnut shows the scores for each of the species. The html version of this plot is interactive, which is useful to obtain the species represented in each segment and the numerical value of the score. As the pairwise structure has the additional column n for the number of observations in each score calculation, the bullseye plot has slices in proportion to n. In this display we see there are far fewer Chinstraps than Gentoos or Adelies. Generally, it is of interest if there is a high difference among the grouped and ungrouped scores. Here the variable bill_dep has a negative overall correlation with each of flip_len, body_mass, bill_len but positive correlation within each of the species groups. The same pattern (known as Simpson’s paradox) is present for the pair bill_len and bill_dep. There are also a number of wedges coloured in grey involving the variable island, representing score values of NA. This is because two of the species (Gentoo and Chinstrap) are located on one island apiece only. For these species, island is constant and so the association value cannot be calculated and is therefore recorded as NA.
@@
To demonstrate the linear display available in bullseye, we switch to a dataset with more variables. We use the acs12 dataset from the package openintro (openintro) which contains results from the US Census American Community Survey in 2012. The dataset has 13 variables, only four of which (income, hrs_work, age, time_to_work) are numeric, the rest are factors. The code segment below calculates ace, cancor and nmi scores. (These three measures were chosen because they are applicable to both numerical and factor variables.)
@@
A display similar to Figure 2 could be constructed with the code
sc_multi |>
 plot_pairwise(interactive=TRUE)
However, rather than displaying all ${{13}\choose{2}}=78$ variable pairs we chose instead to filter out the interesting variable pairs from sc_multi. In this example, we show variable pairs whose maximum score across the three measures exceeds 0.25. The filtered pairwise structure no longer has all (x,y) pairs, so we use the linear display as shown in Figure 4 instead of the matrix display.
@@
  plot_pairwise_linear()
@@
In Figure 4 plot_pairwise_linear uses the default argument geom="point", the alternative geom="tile" shows a filled tile representing the score value and is generally less useful. The function also uses the default argument pair_order = "seriate_max" which sorts variable pairs from highest to lowest maximum score. The alternative pair_order = "seriate_max_diff" sorts pairs by maximum difference, see Figure 5 for an example.
In Figure 4 we see that the mutual information measure (nmi) is consistently less than the ace score. By definition the ace score cannot be less than the cancor score, and they are identical for categorical pairs of variables, for example, (citizen, lang) and (lang, race). The ace scores for income with other variables is much higher than the cancor score, indicating non-linear association. For example, an investigation of the (age, income) scatterplot shows that income goes up with age until about age 40 and then drops off.
@@
pairwise_scores(acs12, by = "race") |>
  mutate(valrange = diff(range(value)),valmax = max(abs(value)), .by=c(x,y))|>
@@
  plot(type="linear", geom="point", pair_order = "seriate_max_diff")
@@ 5  Bullseye and other packages
Finally, the pairwise tibble is easily manipulated to construct general plots: The example code below and resulting display in Figure 6 shows that pivoting the sc_multi tibble containing scores calculated on the acs12 data, gives a dataset where the relationship between the scores can easily be compared using a standard ggplot2 visualisation. Both the scores ace and nmi are designed to capture non-linear associations. Here we see that for these data there is an increasing trend in the relationship between ace and nmi scores, though ace scores are higher. In general, the standardised front-end for calculating and managing pairwise association measures provided by bullseye facilitates comparison of the performance of those measures.
@@ 6  Discussion
In this paper we described a new R package bullseye which fills a gap in providing a standard interface for the calculation of (grouped) pairwise variable scores, a tidy data structure for their manipulation, and a new, optionally interactive, visualisation for the display of multiple pairwise scores. As demonstrated in Section 4.2 the visualisation is particularly useful for easy identification of the presence of Simpson’s paradox. A second linear display is useful to hone in on a filtered set of interesting pairwise scores.
As explained in Section 4.1, the pairwise constructor assumes that pairwise variable scores are symmetric, and similarly at present the visualisation functions assume symmetry. It would however be straightforward to extend to asymmetric measures and visualisations, and furthermore to include scores and visualisations for single variables.
Of course, there are a plethora of methods for measuring pairwise patterns which are not included in our package. For example, tjostheim provides a survey of recent developments in measures of dependence for numerical variables. Recently, lancaster described an alternative to ace for measuring maximal correlation, which is implemented in package lancor (lancor). Salahub present an association measure based on recursive random binning, and provide an associated package AssocBin (assocbin). However, as discussed in Section 5, bullseye visualisations are easily accessible for pairwise scores implemented in R beyond those implemented in our package.
The interactive capabilities of bullseye are currently limited to providing informative tooltips. In future, our package could be used as vehicle for interactively displaying interesting pairwise variable plots. Possibilities in this area have been explored by grimm2017kennzahlenbasierte, who in the archived package mbgraphic, used correlation displays and scagnostics to identify informative scatterplot matrices and parallel coordinate plots embedded in an interactive Shiny application, and also by scagexplorer, though their work is not implemented in an R package.
@@ 7  Acknowledgement
This publication has emanated from research supported in part by a grant from Taighde Eireann – Research Ireland under Grant number 18/CRT/6049. For the purpose of Open Access, the author has applied a CC BY public copyright licence to any Author Accepted Manuscript version arising from this submission
Amit Chinwan
Maynooth University
Hamilton Institute
Maynooth University, Ireland
amit.chinwan.2019@mumail.ie
Catherine B. Hurley
Maynooth University
Department of Mathematics and Statistics
Maynooth University, Ireland
ORCiD: 0000-1721-1511-1101
catherine.hurley@mu.ie